If n(AUB) = 43. n(ANB)=22, and n(B) = 37, find n(A). NA) Enter your answer in the answer box. ?
Answer #1

We know that n(AUB) = n(A) +n(B) -n(AB) .

Therefore, 43= n(A)+37-22 = n(A)+15.

Hence n(A) = 43-15 = 28.